OpenStudy (anonymous):

simplify x/(x^2-y^2)-y/(x^2-y^2)

OpenStudy (anonymous):

= ( x - y ) / (x-y) ( x+y) Can you continue?

OpenStudy (anonymous):

i would cancel the x-y leaving me with 1/x+y right?

OpenStudy (anonymous):

how did u do that because i have several subtration ones i dont understand

OpenStudy (anonymous):

Memorize the essential formula: x² - y² = ( x - y) ( x + y)

OpenStudy (anonymous):

ok so as long as the bottom is the same i can subtract them and then factor it out right

OpenStudy (anonymous):

That's the ultimate basic rule as 2/2 - 1/2 = 1/2

OpenStudy (anonymous):

ok thank you so much for helping me and i think it makes sense now to me on how you were able to get the answer

OpenStudy (anonymous):

I'm glad you ask, otherwise I don't know where to explain :)

OpenStudy (anonymous):

is it x ^6 - y ^ 7